Determine the point-slope form of the line and the slope-intercept form of the line with x-intercept = 4 and y-intercept = -2

Answer by Alan3354(69443)

Slope m = diffy/diffx = 2/4 = 1/2
y-intercept = -2
--> y = (1/2)x - 2